OpenStudy (anonymous):

Nathan Alexander wants to purchase ears of corn for a family reunion picnic. A farmers’ market near his home offers six ears of corn for $1.25, 12 ears of corn for $2.64, or 36 ears of corn for $7.20. Which is the best buy?

Directrix (directrix):

Try getting the price per ear of corn and comparing. 1.25/6 = ? 2.64/12 = ? 7.20/36 = ? @Emilyh117 After you do the divisions, post what you get and we will compare answers, okay?